OpenStudy (anonymous):

please help i'm struggling determine whether the graphs of the pair of lines are parallel x+6=y y-x=-2 select correct choice and if necessary fill in the answer box to complete your choice. the slope of the line x+6=y is ----? the slope of the line is underfined the slope of the line y-x=-2 is---? the slope of the line is uderfined are the graphs of the given equations parallel?

OpenStudy (anonymous):

Solve for y. y=mx+b Compare m

OpenStudy (anonymous):

chaguanas could you be a little more specific please

OpenStudy (anonymous):

in the first equation , x+6 =y y=x+6 the slope of this equation is 1. in the second equation , y-x=-2 y=x-2 the slope of this equation is also1 since the slopes are equal ,the lines are parallel. hope this helps.